OpenStudy (anonymous):

If 5 cm3 of gold has a weight of 96.5 grams, what is the density of gold in grams per cubic centimeter?

OpenStudy (anonymous):

This is the reverse of a question down this evenings list..! use density = mass / volume (and bear in mind that 96.5 grams is mass not weight).. what do you get?

OpenStudy (anonymous):

\[\rho = \frac{ m }{ V }\] where m is gold's mass.., and V is gold's volume..., so: \[\rho = \frac{ m }{ V }\] \[\rho = \frac{ m }{ V }\] \[\rho = \frac{ 96.5 }{ 5 } \] in unit of gr/cm3